QUESTION 1 (C10-1)

Cocao Excavating Inc. has decided to purchase 20 electric vehicles on April 30, 2019, spending

1.75 million plus 13 percent sales tax with a rebate of 2% sales instant on the spot tax break. The company expects to use the vehicles for five years and then sell them for $320,000. Cocao expects the following vehicle use over each year ended April 30:

2019

2020

2021

2022

2023

Km per/yr.

25,000

30,000

20,000

22,000

7,000

To finance the purchase Cocao Excavating Inc. signed a five-year promissory note on December 31, 2017, for 1.4 million with interest paid annual at the market interest rate of 7 percent. The note carries loan covenants that require Cocao to maintain a minimum time interest earned ratio of 3.0 and minimum fixed asset turnover ratio of 1.0. Cocao forecasts that the company will generate the following sales and preliminary earnings (prior to recording depreciation on the vehicles and interest on the note). Income taxes can be ignored

(in thousands)

2019

2020

2021

2022

2023

Revenue

$3,000

$2,750

$3,100

$3,105

3,250

Income before depreciation and interest expense

$1,250

$1,750

$1,900

$2,050

$1,975

REQUIRED:

  1. Calculate the amount of interest expense that would be recorded each year.
  2. Calculate the depreciation expense that would be recorded each year, using the:
    1. Straight-line method
    2. Double Declining method
    3. Units of Production method
  1. Use your answers from Question 1, and 2 to determine net income at the two loan covenant ratios in each year, assuming the company chooses the:
    1. Straight-line method
    2. Double Declining method
    3. Units of Production method
  1. Using your answer to Question 3, indicate whether the loan covenants would be violated under the:
    1. Straight-line method
    2. Double Declining method
    3. Units of Production method
Answer #1
Answer 1
Interest expense (1400000*7%)            98,000
Answer 2 (a) Straight-line depreciation
Cost of assets (1750000*1.13 = 1977500) (1977500-(1977500*2%))      1,937,950
Salvage value          320,000
Depreciable cost      1,617,950
Divided: useful life in Years                       5
Depreciation expense per year          323,590
Answer 2 (b) Double-declining-balance
Depreciation rate under (1/5 year =20%) (20%*2=40%) 40%
Book value of machine does not less than salvage value. For year 4 = 418597 - 320000 salvage value = 98597
Year Beginning balance Depreciation @ 40% ending balance
1          1,937,950           775,180      1,162,770
2          1,162,770           465,108          697,662
3              697,662           279,065          418,597
4              418,597             98,597          320,000
5              320,000                      -            320,000
Answer 2 (C) Units-of-production
Cost of assets (1750000*1.13 = 1977500) (1977500-(1977500*2%))          1,937,950
Salvage value              320,000
Depreciable cost          1,617,950
Divided: useful life in miles Total of Miles (25000+30000+20000+22000+7000)              104,000
Depreciation expense per mile $     15.557212
Year Miles Depreciation expense per mile Depreciation expense
1                25,000 $ 15.557212          388,930
2                30,000 $ 15.557212          466,716
3                20,000 $ 15.557212          311,144
4                22,000 $ 15.557212          342,259
5                   7,000 $ 15.557212          108,900

? X GS HOMEINSERT 2019 10 01.xlsx - Microsoft Excel DEVELOPER - 2 Sign in FILE PAGE LAYOUT FORMULAS DATA REVIEW VIEW P1 : fic? X GS HOMEINSERT 2019 10 01.xlsx - Microsoft Excel DEVELOPER - 2 Sign in FILE PAGE LAYOUT FORMULAS DATA REVIEW VIEW P47 : fo? X GS HOMEINSERT 2019 10 01.xlsx - Microsoft Excel DEVELOPER - 2 Sign in FILE PAGE LAYOUT FORMULAS DATA REVIEW VIEW P71 x fi